OpenStudy (anonymous):

What is the distance between the points (2, 5) and (5, 7)?

OpenStudy (anonymous):

So you want the diagonal distance between the two points. So using pythagoreas, you want the x and y distances between both points. Which would be 3 and 2 3^2+2^2=c^2. So c=sqrt 13

OpenStudy (anonymous):

a) √10 b) √13 c) 5

OpenStudy (anonymous):

\[\sqrt{(5-2)^2+(7-5)^2}=\sqrt{9+4}=\sqrt{13}\]